OpenStudy (jessicawade):

Check my answers? will medal graphs of rational functions 3 questions

OpenStudy (jessicawade):

I need help with these 3 problems. 9. for the function f(x)=3x^2-5x+1 over x +1, identify the slant asymptote A)The slant asymptote for this function is the line y = 3x + 8. ~~~ b)The slant asymptote for this function is the line y = −7(x + 1). c)The slant asymptote for this function is the line y = 3x − 5. D)The slant asymptote for this function is the line y = x + 1. 11. Using the famous Distance equals Rate times Time equation write a specific equation to describe traveling down the road at 65 miles per hour in a car. A) D=65 * T B)65= D*T C) D=T/ 65 D) D= 65/T 14.Suppose that the time it takes to do a job is inversely proportional to the number of workers. That is, the more workers on the job the less time required to complete the job. If it takes 3 workers 16 days to finish a job, how long will it take 8 workers? A)6 days B)14 days C)2 days D)10 days

OpenStudy (jessicawade):

I got A for 11 and D for 14

OpenStudy (jessicawade):

for number 9 my calculator said it was 3x-8 but it wasnt a choice so i went with A

OpenStudy (dumbcow):

ok you got 11 right

OpenStudy (dumbcow):

what calculator are you using? you need to do long division for 9

OpenStudy (jessicawade):

an online one

OpenStudy (dumbcow):

oh ok that works ... yeah its right the slant asymptote is 3x-8

OpenStudy (dumbcow):

maybe theres a typo

OpenStudy (jessicawade):

hmm ok

OpenStudy (jessicawade):

and i think 14 is D

OpenStudy (jessicawade):

@dumbcow